OpenStudy (anonymous):

Someone please help , Franco is traveling to a vacation destination 715 miles from home . on the first day of his trip he travels 390 miles in 6 hours . On the second day of his trip he leaves at 8:00 a.m. and arrives at his destination at 1:00 p.m. What is theunit rate of the total number of miles per hour traveled ? What is the unit rate of of the total number of miles per hour traveled? Convert the information to coordinate points and use the formula y^2-y^1 / x^2-x^1 to answer the question.

OpenStudy (anonymous):

y^2-x-y/x2

OpenStudy (anonymous):

Huh ?

OpenStudy (mertsj):

First Day: 390 miles in 6 hours (6,390) 390/6 mph= 65mph Second Day: 325 miles in 5 hours (5,325) 325/5 mph=65 mph (390-325)/(6-5)=65/1

OpenStudy (anonymous):

thank you .

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

i dont know